OpenStudy (anonymous):

f(x) = 2x-9 , g(x) = 9x-5 ... find (f-g)(x)

OpenStudy (anonymous):

(f-g)(x)= f(x)-g(x)= 2x-9-(9x-5)...

OpenStudy (anonymous):

do you need the answer, or is that good enough?

OpenStudy (anonymous):

so it would be -7x-4 ?

OpenStudy (anonymous):

that's what i got.

OpenStudy (anonymous):

do you have a way to double check the answer?

OpenStudy (anonymous):

no. not that i am aware.

OpenStudy (anonymous):

okay well i'm pretty confident in that. If you get a chance to double check it go ahead, but looks quite correct to me. hope it works out for you.

OpenStudy (anonymous):

I'll let you know if anything convinces me otherwise.

OpenStudy (anonymous):

haha. ok thanks

OpenStudy (anonymous):

yeah, sure thing :D